# One Shot Dominance: Knowledge Poisoning Attack on Retrieval-Augmented Generation Systems

### v1

@@ Abstract
Large Language Models (LLMs) enhanced with Retrieval-Augmented Generation (RAG) have shown improved performance in generating accurate responses. However, the dependence on external knowledge bases introduces potential security vulnerabilities, particularly when these knowledge bases are publicly accessible and modifiable. Poisoning attacks on knowledge bases for RAG systems face two fundamental challenges: the injected malicious content must compete with multiple authentic documents retrieved by the retriever, and LLMs tend to trust retrieved information that aligns with their internal memorized knowledge. Previous works attempt to address these challenges by injecting multiple malicious documents, but such saturation attacks are easily detectable and impractical in real-world scenarios. To enable the effective single document poisoning attack, we propose AuthChain, a novel knowledge poisoning attack method that leverages Chain-of-Evidence theory and authority effect to craft more convincing poisoned documents. AuthChain generates poisoned content that establishes strong evidence chains and incorporates authoritative statements, effectively overcoming the interference from both authentic documents and LLMs’ internal knowledge. Extensive experiments across six popular LLMs demonstrate that AuthChain achieves significantly higher attack success rates while maintaining superior stealthiness against RAG defense mechanisms compared to state-of-the-art baselines.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and found widespread applications in daily life. However, despite their impressive abilities, LLMs still face challenges such as outdated knowledge and hallucination issues as knowledge continues to evolve Achiam et al. (2023); Touvron et al. (2023); Anil et al. (2023). To address these limitations, researchers have introduced the Retrieval-Augmented Generation (RAG) approach to enhance the accuracy of LLM responses Tu et al. (2024); Zhao et al. (2024). RAG consists of three main components: external knowledge base, retriever, and the LLM. Typically, when a user question is received, the retriever searches the external knowledge base for the most relevant information, which is then provided as context to the LLM, enabling it to generate more precise and accurate responses. This technology has been widely adopted in both industrial applications and academic research Al Ghadban et al. (2023); Wang et al. (2024a); Loukas et al. (2023); Kumar et al. (2023); Prince et al. (2024).
@@
While RAG significantly enhances LLM response performance, it also introduces potential security vulnerabilities. Recent studies have demonstrated various attacks targeting either retrievers or LLMs: some attacks manipulate retrievers to fetch maliciously crafted documents, while others exploit backdoors in LLMs to control their responses Tan et al. (2024); Xue et al. (2024); Chaudhari et al. (2024). However, such attacks depend on direct access to RAG pipeline components (retrievers or LLMs), which is typically infeasible in commercial systems where internal implementations are heavily protected. The knowledge base, a more accessible component of the RAG framework, presents a more feasible target for attacks. For instance, when users question about current events, RAG systems typically search Wikipedia Thakur et al. (2021) as their knowledge base to generate responses Chen et al. (2024). Since Wikipedia allows users to modify and upload content, its knowledge base is particularly vulnerable to manipulation Carlini et al. (2024).
Poisoning knowledge bases faces significant challenges, as the effectiveness of injected malicious content is heavily influenced by two aspects: (1) the injected content needs to compete with multiple authentic documents retrieved by the system, where these documents often present consistent factual information that contradicts the injected content Xie et al. (2023); (2) LLMs tend to prefer retrieved information that matches their internal memorized knowledge from pre-training Jin et al. (2024). These factors create strong resistance against malicious content manipulation. Previous work like PoisonedRAG Zou et al. (2024) attempts to overcome these challenges by injecting multiple targeted documents for a single question, ensuring all retrieved documents are poisoned ones. While this saturation attack achieves high attack success rates, it is easily detectable due to the unusual pattern of multiple similar malicious documents appearing in the retrieval results. When limiting to a single poisoned document per question to avoid detection, the aforementioned challenges become even more pronounced. The example is shown in Figure 1.
To tackle the first challenge, we draw inspiration from the Chain-of-Evidence (CoE) theory in criminal procedure law, which requires decisive evidence in judicial decisions to demonstrate both relevance (relation to the case) and interconnectedness (mutual support among evidence) Murphy (2013). Drawing parallels to how LLMs select correct answers from external knowledge, we construct targeted documents as evidence chains for specific questions (incorporating the core intent of the question and establishing connections between keywords), making them more appealing to LLMs compared to other retrieved knowledge. To address the second challenge, inspired by the authority effect in social psychology Cialdini and Goldstein (2004), we craft statements that include recent timelines and endorsements from authoritative institutions for target answers, thereby strengthening the LLM’s reliance on external knowledge and disrupting the influence of its internal knowledge. Based on these insights, we propose AuthChain, a novel single document poisoning attack method that combines CoE and authoritative statements.
@@
We evaluate AuthChain across six popular LLMs, demonstrating significantly higher attack success rates than the state-of-the-art baselines. Furthermore, under RAG defense frameworks, AuthChain exhibits superior stealthiness by more effectively evading detection mechanisms. The reproduction package is available at: https://anonymous.4open.science/r/AuthChain-45E8.
@@ 2.2 Existing Attacks on RAG
Traditional machine learning attacks have primarily focused on white box scenarios, where attackers can directly compromise model training through data poisoning and backdoor attacks Carlini et al. (2024); Biggio et al. (2012); Liu et al. (2018). In RAG systems, several sophisticated white box approaches have been developed. Jamming optimizes token selection process and introduces instruction attacks to alter LLM’s fundamental behavior Shafran et al. (2024). Other researchers have explored trigger-based attacks: Phantom introduces trigger-specific malicious behaviors Chaudhari et al. (2024), LIAR improves attack success by alternating between retriever and generator targeting Tan et al. (2024), and BadRAG enables flexible trigger selection for privacy compromise and denial of service Xue et al. (2024).
However, these white box approaches become impractical when targeting commercial RAG systems where the LLM and retriever are managed by major tech companies Gu et al. (2017); Shafahi et al. (2018). This limitation has led to the emergence of black box attacks that target the knowledge database: a more accessible attack surface. In the black box setting, PoisonedRAG Zou et al. (2024) constructs poisoned documents by combining the original questions with LLM-generated supporting content. The original questions are used as prefixes to ensure high retrieval probability, while the LLM is prompted to generate fake yet plausible knowledge that guides the model towards the attacker’s desired outputs. Meanwhile, HijackRAG Zhang et al. (2024) employs a three-component strategy combining retrieval text (original questions), hijack text (selected from public prompt hacking datasets), and instruction text (explicitly guiding the model to generate desired responses). However, these black-box methods require inserting multiple poisoned samples for a single target question, which increases the risk of being detected by RAG systems. In this work, we explore the attack method that only require inserting a single poisoned document for each target question in the black-box setting.
@@ 3 Methodology
To address the challenges from both authentic documents and LLMs’ internal knowledge, we draw inspiration from two established theories in different domains. First, to help the injected content stand out among multiple authentic documents, we leverage the Chain-of-Evidence (CoE) theory from criminal procedure law. This theory emphasizes that decisive evidence must demonstrate both relevance to the case and interconnectedness among different pieces of evidence to establish credibility. Similarly, by constructing poisoned content as interconnected evidence chains that closely relate to the question’s key elements, we can make it more convincing to LLMs compared to other retrieved documents. Second, to overcome LLMs’ tendency to trust information aligned with their internal knowledge, we draw upon the authority effect from social psychology, which suggests that people are more likely to accept information from authoritative sources even when it conflicts with their existing beliefs.
Based on these insights, we propose AuthChain, as shown in Figure 2, which consists of three main stages: (1) Question-Based Hint Generation analyzes the input question by extracting CoE features and transforming them into structured hints, which serve as the foundation for subsequent content generation. (2) CoE Content Generation employs an iterative feedback-enhanced approach to generate CoE content that fully covers the extracted CoE features, ensuring such knowledge can be retrieved at top ranks by retrievers and preferred by LLMs among conflicting external documents. (3) Authority Content Generation builds upon the generated hints to create authoritative content with domain-specific details, designed to overcome LLM’s reliance on internal knowledge while enhancing the credibility of targeted answers.
3.1 Question-Based Hint Generation
To construct external knowledge that can be effectively retrieved by the retriever and preferred by LLMs, we aim to characterize the essential features of such knowledge. Drawing inspiration from the law of criminal procedure, where judicial decisions require a Chain-of-Evidence (CoE) through evidence collection Edmond and Roach (2011); Murphy (2013), we identify two key properties that make knowledge compelling: relevance (closely related to the question) and interconnectivity (evidence nodes mutually supporting each other). We analogize judicial decisions to the scenario in which LLMs identify answers from external knowledge in response to user questions.
To construct external knowledge that forms a CoE, which we hypothesize is preferred by LLMs, we extract three key CoE features from the targeted question:
@@ 3.1 Question-Based Hint Generation
  • Intent extracted as a noun or noun phrase from the question, represents the user’s ultimate goal or desired answer. This helps ensure alignment between the question’s purpose and the knowledge to be generated.
  • Evidence Nodes are key entities identified from the question that form essential stepping stones in the reasoning path from the question to its answer. It ensures the consistency between external knowledge and the question from the entity perspective.
  • Evidence Relations capture the logical connections between evidence nodes as expressed in the question. These relations help preserve the question’s inherent reasoning structure in the generated knowledge.
The example of CoE and CoE features are presented in Appendix A.
@@
These CoE features serve as crucial guidance for subsequent knowledge generation, ensuring the generated content maintains strong relevance to the original question while preserving coherent internal logic. To effectively extract these CoE features, AuthChain leverage the LLM with an enhanced prompting strategy. Building upon the prompt template from  Li et al. (2023), the LLM incorporate few-shot learning by including 5 carefully selected input-output examples, which significantly improves the feature extraction performance. The detailed prompt template and examples are provided in Appendix C.
Following the CoE feature extraction, AuthChain leverage a content agent to generate a question-based hint, which serves as a foundation for subsequent knowledge generation, ensuring the generated content maintains strong semantic relevance to the targeted question. Specifically, we feed the content agent with three inputs: the extracted CoE features, the targeted question, and its targeted answer. Through carefully designed prompts, we guide the agent to generate the question-based hint that seamlessly incorporates CoE features while maintaining semantic coherence. The detailed prompt template for the content agent is provided in Appendix D.
@@ 3.2 CoE Content Generation
To generate content with comprehensive CoE features that support the target answer, we implement an iterative refinement process.
Initially, both the question-based hint and the extracted CoE features are fed into a CoE judge agent. The CoE judge agent evaluates whether the input content fully incorporates all CoE features. If complete coverage is confirmed, the content is directly output. When certain features are missing, the judge agent provides specific advises for incorporating these missing CoE features, such as enhancing evidence nodes or clarifying evidence relations. These advises, together with the current content, are then forwarded to a revise agent for content refinement. This iterative evaluation and revision process continues until the judge agent confirms complete coverage of all CoE features, at which point the current content is finalized as the CoE content. The detailed prompt templates for both the CoE judge agent and revise agent are provided in Appendix E.
@@ 3.3 Authority Content Generation
Considering that LLM’s internal knowledge could interfere with the control of poisoned content over LLM outputs, we aim to reduce LLM’s reliance on internal knowledge and encourage it to rely more on external knowledge when responding to target questions. Drawing inspiration from the authority effect in social psychology Cialdini and Goldstein (2004), we hypothesize that content endorsed by authoritative institutions, coupled with recent timeline statements, can effectively redirect LLM’s attention toward external information while minimizing interference from its internal memory. Specifically, AuthChain feed the question-based hint, targeted answer, and CoE features into an authority generator agent. The inclusion of CoE features aims to enhance the content’s likelihood of being retrieved within the top-k results by increasing its semantic similarity with the question through the incorporation of evidence nodes from CoE features. The agent analyzes the hint’s context to identify the most suitable authoritative institution for endorsement. It then synthesizes this institutional backing with the latest timeline to validate the target answer, ultimately producing authority content. Appendix F shows the prompt for the authority generator agent.
Finally, AuthChain integrate the CoE content with the authority content to create the final poisoned document, which is then injected into the knowledge base. The example of poisoned document generated by AuthChain is shown in Appendix B.
@@ 4.1 Research Questions
RQ1: How effective are the single poisoned document generated by AuthChain at manipulating the output of various LLMs?
RQ2: How effective is AuthChain at evading detection under RAG-based defense frameworks?
RQ3: Can authority content overcome LLMs’ memorized biases towards external knowledge, while CoE maintains dominance across conflicting external sources?
@@ 4.2 Dataset
In our experiments, we utilize three information retrieval datasets from the Beir benchmark Thakur et al. (2021): Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), and MS-MARCO Nguyen et al. (2016). Each dataset contains a knowledge database and question-answer pairs. For NQ and HotpotQA, the knowledge databases are derived from Wikipedia, containing 2,681,468 and 5,233,329 texts respectively. MS-MARCO’s knowledge database is collected from web documents through Bing search engine Microsoft , comprising 8,841,823 texts. Following the evaluation setup of Zou et al. (2024), we use their constructed test set containing 100 query-answer pairs for each dataset.
4.3 RAG Setup
The RAG implementation consists of three components: knowledge base, retriever, and LLM. We use the original knowledge bases provided by the datasets, Contriever Izacard et al. (2021) as the retriever, and six different LLMs including three open-sourced LLMs (GPT-3.5 OpenAI (2022), GPT-4 Achiam et al. (2023), GPT-4o OpenAI (2024)) and three closed-sourced LLMs (LLama3-8B Touvron et al. (2024), LLama3-70B Touvron et al. (2023), Deepseek-V3-0324 DeepSeek (2024)). For each query, the retriever selects the top-5 most relevant documents as context for the LLM. For all LLMs, we use the system prompt from Zou et al. (2024) to instruct models to generate answers based on the retrieved documents.

### v2

@@ Abstract
Large Language Models (LLMs) enhanced with Retrieval-Augmented Generation (RAG) have shown improved performance in generating accurate responses. However, the dependence on external knowledge bases introduces potential security vulnerabilities, particularly when these knowledge bases are publicly accessible and modifiable. While previous studies have exposed knowledge poisoning risks in RAG systems, existing attack methods suffer from critical limitations: they either require injecting multiple poisoned documents (resulting in poor stealthiness) or can only function effectively on simplistic queries (limiting real-world applicability). This paper reveals a more realistic knowledge poisoning attack against RAG systems that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions involving complex relationships between multiple elements. Our proposed AuthChain address three challenges to ensure the poisoned documents are reliably retrieved and trusted by the LLM, even against large knowledge bases and LLM’s own knowledge. Extensive experiments across six popular LLMs demonstrate that AuthChain achieves significantly higher attack success rates while maintaining superior stealthiness against RAG defense mechanisms compared to state-of-the-art baselines.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and found widespread applications in daily life. However, despite their impressive abilities, LLMs still face challenges such as outdated knowledge and hallucination issues as knowledge continues to evolve Achiam et al. (2023); Touvron et al. (2023); Anil et al. (2023). To address these limitations, researchers have introduced the Retrieval-Augmented Generation (RAG) approach to improve LLMs Tu et al. (2024); Zhao et al. (2024). This technology has been widely adopted in both industrial applications and academic research Al Ghadban et al. (2023); Wang et al. (2024a); Loukas et al. (2023); Kumar et al. (2023); Prince et al. (2024).
@@
Although RAG significantly improves the performance of the LLM response, it introduces potential security vulnerabilities. The security of RAG systems is influenced by both the inherent safety mechanisms of LLMs Tan et al. (2024); Xue et al. (2024); Chaudhari et al. (2024) and the potential manipulation of external context Zou et al. (2024); Zhang et al. (2024). Attackers can attempt to mislead LLMs into generating incorrect outputs by poisoning knowledge bases with malicious content. This attack surface is particularly concerning as knowledge bases are often the most accessible component of RAG systems. For example, when RAG systems utilize public resources like Wikipedia Thakur et al. (2021) for current events information Chen et al. (2024), attackers could exploit the open editing nature of these knowledge bases to inject malicious content that may alter the LLM’s outputs.
While previous studies have exposed knowledge poisoning risks in RAG systems, existing attack methods suffer from critical limitations: they either require injecting multiple poisoned documents (resulting in poor stealthiness) or can only function effectively on simplistic queries (limiting real-world applicability) Zou et al. (2024); Zhang et al. (2024).
This paper reveals a more realistic knowledge poisoning attack against RAG systems that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions involving complex relationships between multiple elements. As shown in Figure 1, such single document poisoning faces significant challenges. When the malicious attacker constructs a targeted document replacing “Kosovo” with “Nagorno-Karabakh”, the attack fails because: 1) The LLM prefers answers that appear across multiple retrieved documents rather than from a single source, as indicated by the multiple documents showing repeated mentions of “Kosovo”; 2) The LLM favors answers that align with its internal knowledge - in this case, its embedded knowledge correctly states that “Kosovo’s declaration did not violate international law”; 3) Simply injecting the question context into a single document creates unnatural content patterns that reduce the document’s credibility during retrieval and reasoning. These challenges make single document poisoning particularly difficult, as the attack must overcome both the LLM’s preference for consensus across multiple documents and alignment with its internal parameterized knowledge.
To address these challenges, we propose AuthChain, a novel single document poisoning attack method that stealthily executes knowledge poisoning in RAG scenarios. Our approach follows three key principles for crafting an effective malicious document, with each progressively strengthens the attack. First, ensuring visibility. The poisoned document must stand out among vast external information sources. We achieve this through precise alignment with the question’s underlying objective. When external knowledge perfectly mirrors the core intent of a query, retrievers naturally prioritize it in their rankings, while LLMs tend to focus more on such intent-aligned information during their reasoning process. Second, guaranteeing retrievability and competitiveness. Even for complex queries involving multiple elements with complex relations, the document must remain a top candidate and outperform other retrieved knowledge. We accomplish this by structuring it as a self-contained evidence chain, preserving all key question elements and their logical relationships. This complete chain structure not only boosts retrieval rankings but also makes our content more compelling than fragmented knowledge pieces that only partially match the question’s logic. Third, overcoming LLMs’ internal knowledge bias. When LLMs consider multiple information sources, they tend to favor external knowledge that aligns with their internal knowledge. To counteract this bias, we strategically incorporate authority signals into our document, such as endorsements from authoritative institutions and recent timestamps. These authority signals help position our content as a more current and authoritative source compared to LLMs’ static internal knowledge, effectively overcoming their inherent preference for internally knowledge.
@@
We evaluate AuthChain across six popular LLMs, achieving 21.7%-46.5% improvements in attack success rates compared to state-of-the-art baselines. Furthermore, under two RAG defense frameworks, AuthChain exhibits superior stealthiness by more effectively evading detection mechanisms. The reproduction package is available at: https://anonymous.4open.science/r/AuthChain-45E8.
@@ 2.2 Existing Attacks on RAG
In RAG systems, several white box approaches have been developed. Jamming optimizes the token selection process and introduces instruction attacks to alter LLM’s fundamental behavior Shafran et al. (2024). Other researchers have explored trigger-based attacks: Phantom introduces trigger-specific malicious behaviors Chaudhari et al. (2024), LIAR improves attack success by alternating between retriever and generator targeting Tan et al. (2024), and BadRAG enables flexible trigger selection for privacy compromise and denial of service Xue et al. (2024).
However, these white box approaches become impractical when targeting commercial RAG systems where the LLM and retriever are managed by major tech companies Gu et al. (2017); Shafahi et al. (2018). This has led to black box attacks that target the knowledge database as a more accessible attack surface. Recent works like PoisonedRAG Zou et al. (2024) and HijackRAG Zhang et al. (2024) propose methods combining original questions with manipulated content to achieve attacks. However, these methods face critical limitations in real-world scenarios: they require injecting multiple poisoned documents, resulting in poor stealthiness, and can only function effectively on simplistic queries. In this work, we explore a more realistic attack method that only requires inserting a single poisoned document while remaining effective for complex multi-hop questions involving intricate relationships between multiple elements.
@@ 3 Methodology
To achieve effective single document poisoning in RAG, we propose AuthChain. Unlike previous approaches that require multiple poisoned documents or only effectively work for simple queries, AuthChain aims to create a malicious document that can not only be successfully retrieved among vast external knowledge sources, but also effectively compete with both authentic documents and LLMs’ internal knowledge, even for complex multi-hop questions. Our key insight is that the document’s influence can be progressively enhanced through three aspects: maximizing visibility during retrieval by precisely aligning with the question’s intent, enhancing both retrievability and competitiveness through complete evidence chains that maintain high retrieval rankings while outperforming fragmented external knowledge, and overcoming LLMs’ internal knowledge through authority reinforcement.
As shown in Figure 2, we implement these design principles through three main stages: (1) Intent-Based Content Generation focuses on maximizing document visibility by extracting three key features from the input question (the intent, key elements, and their relationships) and generates intent-based content. (2) CoE Content Generation aims to maintain high retrievability by constructing complete evidence chains. Using the extracted features and intent-based content, it generates Chain-of-Evidence (CoE) content that fully preserves the question’s semantic structure by covering its core objective, all key elements, and their relationships. (3) Authority Content Generation enhances document trustworthiness by incorporating domain-specific authority signals. Building upon the intent-based content, it creates authoritative content by incorporating institutional affiliations and recent timestamps, while maintaining professional formatting consistent with authoritative sources.
3.1 Intent-Based Content Generation
To maximize document visibility during retrieval, AuthChain first extracts key features from the question and generates an intent-based content that guides subsequent content generation. As shown in Figure 2, this stage consists of two main steps: feature extraction and intent-based content generation.
For feature extraction, we systematically analyze the question to capture both its intent and evidence chains:
@@ 3.1 Intent-Based Content Generation
  • Intent, extracted as a noun or noun phrase, represents the question’s ultimate goal. This helps ensure the generated content directly addresses what LLMs prioritize during retrieval and reasoning.
  • Evidence Chains, consisting of evidence nodes and their relations, captures the question’s logical structure: Evidence Nodes are key entities in the question that serve as critical components. Evidence Relations represent logical connections between these nodes.
@@
The example of question-derived features are presented in Appendix A. To effectively extract these features, AuthChain employs an LLM-based extraction approach enhanced with few-shot learning. Building upon the prompt template from  Li et al. (2023), we incorporate 5 carefully selected examples to improve extraction performance. The detailed prompt template and examples are provided in Appendix D.
Given the extracted intent, targeted question and answer, AuthChain prompts an intent agent to generate intent-based content. Such intent-driven content naturally aligns with the question’s essential requirements, which helps it achieve higher retrieval rankings and receive increased attention during LLM reasoning processes. Besides, it guides the subsequent knowledge generation to produce evidence that directly addresses the question’s central concerns. The detailed prompt template for the intent agent is provided in Appendix H.
@@ 3.2 CoE Content Generation
To both maintain high retrievability and outperform other external knowledge sources, AuthChain constructs complete evidence chains that preserve all question elements and their logical connections. While the intent-based content provides initial alignment with the question, we need to ensure the generated content comprehensively covers all extracted evidence nodes and their relationships.
As shown in Figure 2, we implement an iterative refinement process to construct the Chain-of-Evidence (CoE) content. First, both the intent-based content and the extracted evidence chains are input to a CoE judge agent. This agent evaluates whether the content fully incorporates all evidence nodes and their relationships. If complete coverage is confirmed, the content is directly output as CoE content. Otherwise, the judge agent provides specific advice for incorporating missing elements, such as adding absent evidence nodes or establishing semantic relationships between nodes.
These suggestions, along with the current content, are then forwarded to a revise agent for refinement. This iterative evaluation and revision process continues until the CoE judge agent confirms complete preservation of the evidence chains, at which point the current content is finalized as the CoE content. The detailed prompt templates for both agents are provided in Appendix F.
@@ 3.3 Authority Content Generation
To mitigate LLM’s reliance on internal knowledge when answering questions, we need to make external knowledge more compelling and trustworthy. Drawing inspiration from the authority effect in social psychology Cialdini and Goldstein (2004), we hypothesize that content endorsed by authoritative institutions, coupled with recent timeline statements, can effectively redirect LLM’s attention toward external information while reducing reliance on its internal knowledge.
AuthChain employs an authority generator agent that takes the intent-based content, targeted answer, and question-derived features as input. By incorporating these features, the generated content maintains stronger semantic alignment with the original question, facilitating better retrieval. The agent first analyzes the intent-based content context to identify the most suitable authoritative institution for endorsement. It then synthesizes this institutional backing with recent timeline information to validate the targeted answer, ultimately producing authority content. The detailed prompt template for the authority generator agent is provided in Appendix G.
Finally, AuthChain integrates CoE content with authority content to create the final poisoned document, which is then injected into the knowledge base. The example of poisoned document generated by AuthChain is shown in Appendix B.
@@ 4.1 Research Questions
RQ1: How effective is the single poisoned document generated by AuthChain in manipulating the output of various LLMs?
RQ2: How effective is AuthChain in evading detection under RAG-based defense frameworks?
RQ3: Can authority content overcome LLMs’ internal knowledge bias while CoE effectively competes with conflicting documents?
@@ 4.2 Dataset
We evaluate AuthChain on three widely-used question answering datasets in knowledge base poisoning attacks: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), and MS-MARCO Nguyen et al. (2016). These datasets have been consistently used in prior poisoning attacks Zou et al. (2024); Zhang et al. (2024) due to their diverse knowledge sources (Wikipedia and web documents) and different question complexity levels. Following Zou et al. (2024), we use their constructed test set containing 100 query-answer pairs for each dataset to enable direct comparison with baseline attacks.
4.3 Subject RAG System Setting
We use the original knowledge bases provided by the datasets, Contriever Izacard et al. (2021) as the retriever, and six different LLMs, including three open-source LLMs (GPT-3.5 OpenAI (2022), GPT-4 Achiam et al. (2023), GPT-4o OpenAI (2024)) and three closed-source LLMs (LLama3-8B Touvron et al. (2024), LLama3-70B Touvron et al. (2023), Deepseek-V3-0324 DeepSeek (2024)). For each query, the retriever selects the top-5 most relevant documents as the context.